# Unravelling the asphericities in the explosion and multi-faceted circumstellar matter of SN 2023ixf

### v1

@@ 3.3 Imaging Polarimetry and Spectropolarimetry
We obtained spectropolarimetric observations of SN 2023ixf over five epochs on 2023 June 14, 15, 16, 18, and 19 using the CCD Imaging/Spectropolarimeter (SPOL; Schmidt et al., 1992) mounted on the 2.3 m Bok Telescope at Steward Observatory, Kitt Peak, Arizona. We used the 600 l/mm grating with a spectral coverage of 4000-7550Å and obtained a spectral resolution of 26.53 Å using a slit size of 0.4 mm (corresponding to 4.1” in the sky). In addition to this, we used a Hoya L38 blocking filter with a cut-off wavelength of 380nm. We used polarized standards (Hiltner 960; HD 155528) to calibrate the position angle, with the average discrepancy between the observed and expected position angles for multiple standards being less than 1∘ (Schmidt et al., 1992). Furthermore, the unpolarized standard (BD+28°4211) observations confirmed that SPOL exhibited low instrumental polarization, measuring at less than 0.1%. We use the same data acquisition and reduction procedure mentioned in Bilinski et al. (2018). The polarized spectra corresponding to four exposures that sample 16 waveplate positions were flat-field corrected, background subtracted, and flux calibrated using the SPOL polarization data reduction pipeline to obtain the Stokes parameters q, u, p, and polarization angle.
@@ 4 Photometric Evolution
Panchromatic light curves covering ultraviolet, optical, and near-infrared wavelengths from various facilities are shown in Figure 2, displaying the well-sampled SN evolution spanning 1.4 d – 180 d after the explosion. All the magnitudes presented in this paper were calibrated to the AB system using the transformations described in Blanton & Roweis (2007).
@@ 4.1 Rise times, decline rates and plateau-length
SN 2023ixf shows a sharp rise to the peak in the ultraviolet bands - and (i.e. blue peak) with a rise time of 4.5 d. The rise time is 6 d in -band, which is faster than the prototypical Type II SN 1999em (with minimal signatures of interaction), which displayed a rise time of 10 d in -band (Leonard et al., 2002). The steeper rise to the maximum for SN 2023ixf is attributable to the interaction of the SN ejecta with CSM in its immediate vicinity. This was also ascertained from the follow-up of a large sample of infant SNe II by Förster et al. (2018), where the fast-rising light curves were only reproducible via an interaction with a proximal CSM. The steepening of the light curve (and faster rise) is more pronounced in the bluer bands ( 0.5 m) due to heating as discussed by Morozova et al. (2017); Moriya et al. (2018). The light curves in the redder bands () show a steep rise to a distinct shoulder, i.e., red shoulder at 4.5 d and a gradual ascent to the eventual maximum, i.e., red peak at 16 d.
The presence of the red shoulder is driven by the flux excess due to the interaction of SN 2023ixf with a confined CSM since it coincides with the timing of the peak in the UV wavelengths (i.e., blue peak). A similar shoulder and peak is seen in the simulated red-band light curves of SNe II with strong wind models ( 10-3 ) by Dessart et al. (2017). In addition, the rise to the red peak seems to stem from the adiabatic cooling of shock-heated SN ejecta, which results from the migration of the spectral energy distribution (SED) into the redder bands, despite a declining net bolometric luminosity (Singh et al., 2019b). The rise time for the red peak is higher than that of SNe IIP (7.0 0.3 d) and in the ballpark of SNe IIL (13.3 0.6 d) (Gall et al., 2015). This is suggestive of the higher energy/mass ratio in SNe IIL as seen in SN 2023ixf (see Section 7).
Post-maximum, the multi-band light curves of SN 2023ixf settle onto a plateau of roughly 75 d in and other redder bands before transitioning to the radioactively-powered tail phase at 90 d. The plateau length is at the shorter end of the typical plateau length of 100 d for SNe II, hence putting SN 2023ixf amongst some of the rarely observed short plateau SNe (Hiramatsu et al., 2021; Teja et al., 2022). We estimated the plateau decline rates of SN 2023ixf following the prescription of Anderson et al. (2014a). The -band light curve of SN 2023ixf showed an early plateau decline rate (s1) of and a late-plateau decline rate (s2) of . The late-plateau decline rate of SN 2023ixf is much higher than the mean decline rate of 1.3 inferred for SNe II (Anderson et al., 2014a). The tail phase decline rate (s3) of SN 2023ixf is which is faster than the characteristic decline rate of 56Co to 56Fe (i.e., 0.98 ) indicating incomplete trapping of and -rays.
@@ 4.2 Light curve comparisons with other SNe II
The comparison of -band absolute magnitude light curve of SN 2023ixf is shown in the top panel (A) of Figure 3. Owing to the short plateau length and fast declining nature of SN 2023ixf, the comparison sample consists of normal Type IIP SNe: SN 2006bp (Quimby et al., 2007), SN 2007od (Andrews et al., 2010), SN 2013fs (Bullivant et al., 2018), SN 2016gfy (Singh et al., 2019a), 2017eaw (Szalai et al., 2019) and SN 2020tlf (Jacobson-Galán et al., 2022); short-plateau Type IIP SNe: SN 2018gj (Teja et al., 2023b), SN 2020jfo (Teja et al., 2022), and Type IIL SNe: SN 2013by (Valenti et al., 2015), SN 2014G (Terreran et al., 2016), SN 2017ahn (Tartaglia et al., 2021) and SN 2020pni (Terreran et al., 2022). The peak -band luminosity of SN 2023ixf is –18.2 mag, significantly brighter than the average peak luminosity of SNe II (i.e., –16.74 mag) inferred by Anderson et al. (2014a), which is dominated by the population of slow-declining SNe II. The peak luminosity of SN 2023ixf is similar to the SNe IIL, namely SN 2013by, SN 2014G, and SN 2017ahn, but slightly fainter than that of SN 2020pni; however, it is brighter than the majority of normal and short-plateau SNe IIP. The plateau-decline rate of SN 2023ixf is similar to that of SN 2014G ( 1.7 ) and SN 2013by ( 1.5 ). The 75 d plateau length of SN 2023ixf is similar to that of SN 2014G, SN 2013by and SN 2013fs. We infer a 1.8 0.1 mag drop from the plateau phase to the tail phase for SN 2023ixf, similar to that of SN 2014G ( 2.0 mag) and SN 2013by ( 1.8 mag). Overall, SN 2023ixf shows remarkable photometric resemblance in peak-luminosity, plateau decline rates, plateau length, and plateau drop to SN 2013by and SN 2014G.
@@
We further compare SN 2023ixf to SNe II with CSM interaction, followed up extensively in UV by Swift over the last two decades. The data was downloaded from the Swift Optical/Ultraviolet Supernova Archive (SOUSA777https://pbrown801.github.io/SOUSA/, Brown et al., 2014b) and the Vega-mag were transformed to AB-mag for consistency. SN 2023ixf shows a steep rise in the UV fluxes in and bands, which has been seen only in SN 2013fs, SN 2020pni and SN 2020tlf. In the exhaustive sample of SNe II observed by Swift, even though SN 2023ixf is not the earliest followed-up Type IIP/L SNe, the and bands exhibit a rapid surge in the flux, brightening by over 3 mags in 3 days, before reaching a peak magnitude of –20 mag at 4.5 d. This prolonged brightening observed in the initial phase of the UV light curve indicates a shock breakout within a compact and dense CSM, leading to a more luminous and elongated shock breakout event (Ofek et al., 2010). Such a distinct signature in the early UV light curves has been observed in only a limited number of SNe II. The detection of a UV burst extending over 1-day in PS1-13arp by Gezari et al. (2015) was the first observation hinting at the possibility of the shock breaking out into a confined CSM. In our comparison sample, SN 2020pni (Terreran et al., 2022) and SN 2020tlf (Jacobson-Galán et al., 2022) showed a similar intensification of the early phase UV light curve.
@@
The Swift UV light curves of SNe II typically decline rapidly due to the rapid cooling of the SN ejecta and the prevalence of metal absorption features, which eats away at the UV flux as the ejecta cools. We discern the evidence of the same in the FUV spectra of SN 2023ixf obtained by Teja et al. (2023a) at 23 d, which shows an entirely featureless spectrum owing to the line-blanketing from the iron-group elements (Bufano et al., 2009). However, even with severe line blanketing during the early plateau phase, we still infer a decent contribution ( 10%, see Panel B in Figure 4) of UV flux to the overall bolometric light curve (i.e., a UV excess) during the late-plateau phase ( 50 d). This likely suggests continued interaction with CSM despite the absence of discernible signatures in the spectral sequence outlined in Section 5.2. This aligns with the inference from the theoretical models of interacting SNe II by Dessart et al. (2022) that display early interaction frequently display a surplus of UV radiation during late phases.
@@ 4.3 Bolometric Light Curve, Temperature and Radius Evolution
We computed the pseudo-bolometric light curves of SN 2023ixf using data from GIT, Kanata, Iriki, Nayoro, and Swift/UVOT using SuperBol (Nicholl, 2018). The bolometric luminosity, color temperature (), and radius () evolution of the layer of thermalization were estimated using the blackbody fits to the SED of the SN at each epoch. The missing data in certain filters over the intermediate epochs was interpolated using a low-order spline. We computed the bolometric light curve in 3 wavelength bins, i.e., UVOIR (0.16 – 2.35 m), UVO (0.16 – 0.85 m) and OIR (0.38 – 2.35 m) and is shown in Panel A of Figure 4.
Henceforth, we refer to the UVOIR pseudo-bolometric light curve as the bolometric light curve in our discussion. In an RSG, the shock breakout from the surface of the progenitor happens within an hour of the core collapse, followed by a rapid cooling due to the rapid expansion driven by the shock (Falk & Arnett, 1977). This would lead to a rapidly increasing photospheric radius and a decrease in the temperature of the SN ejecta within a couple of hours of the explosion. However, examining the color temperature and radius evolution shown in Panel C of Figure 4, shows a steep increase in the temperature from around 14,000 K to 35,000 K over a duration of 2.2 0.1 d and a relatively-flat radius evolution at (2.0 0.2) cm (13 1 AU). The extended heating at a near-flat radius, in addition to the prolonged brightening in the UV flux, as discussed in Section 4.2, further strengthens that in SN 2023ixf, the shock broke out inside a compact and dense CSM surrounding the progenitor.
From the epoch of the first spectral observations (i.e., 1.1 d) of SN 2023ixf by Perley & Gal-Yam (2023), the luminosity, = (1.28 0.11) , and effective temperature of (14 2) rises ten folds and two folds, respectively, in a span of just 1 d. However, due to the ensued heating, the bolometric light curve of SN 2023ixf flux peaked later on 4.5 d with a luminosity of 2.5 0.3 . The temporal evolution of UV and NIR fraction of the pseudo-bolometric luminosity is shown in Panel B of Figure 4. The contribution of the UV flux (0.16 – 0.38 m) to the bolometric flux stays at roughly about 85% even beyond the epoch of shock breakout until the bolometric maximum (and the UV peak) is reached indicating an ongoing source of heating. The UV flux falls gradually until the end of the plateau phase as the SN enters the photospheric phase powered by recombination. If we ignore the early UV data for SN 2023ixf, the OIR bolometric light curve (see Panel A in Figure 4) underestimates the bolometric luminosity by an order of magnitude, emphasizing the importance of UV observations of infant CCSNe and its importance to detailed hydrodynamical modeling.
@@
Upon recombination of all the hydrogen in the outer envelope, SNe II post the plateau phase transitions to the tail phase powered by the decay of , which thermalizes the SN ejecta through the emission of -rays and . We computed the 56Ni-mass using the relation postulated by Hamuy (2003). The mean tail luminosity of SN 2023ixf around 145 d yields a mass of . Upon comparison with SN 1987A at a similar phase which synthesized a -mass of 0.075 0.005 (Turatto et al., 1998), a -mass of is estimated for SN 2023ixf. The above techniques assumed complete trapping of -rays by the SN ejecta. However, that is not always true, especially for short-plateau SNe due to a thinner hydrogen envelope (Teja et al., 2022) and may lead to underestimation of the yield. Hence, we modeled the late-phase UVOIR bolometric light curve of SN 2023ixf until 150 d using the analytical formulation from Valenti et al. (2008). We derived a total -mass of 0.059 0.001 and a characteristic -ray trapping timescale of 220 3 d, indicating a short-lived trapping and a shallower envelope. We derive this as the upper limit of yield since there might be an additional contribution to the tail phase from CSM interaction as mentioned in Section 4.2.
@@ 5.1 Flash Spectroscopy Phase 8 d
Earlier spectroscopic studies on SN 2023ixf (Yamanaka et al., 2023; Jacobson-Galan et al., 2023; Smith et al., 2023; Bostroem et al., 2023; Teja et al., 2023a) showed a prominently blue continuum coupled with highly ionized features of C IV, N IV, C III, N III and He II along with Balmer features beginning 1.1 d from the estimated date of explosion. The features suggest flash-ionization and recombination of the confined dense CSM engulfing the progenitor of SN 2023ixf. Notably, these studies observed an increase in the relative line strengths of the highly ionized species such as N IV and C IV in contrast to N III and C III, respectively, implying a delayed ionization of the CSM. The increase in line fluxes of N IV, C IV and He II noticed in the spectral sequence of SN 2023ixf spanned 4 d from the explosion (Smith et al., 2023), which roughly corresponds to the peak of UV light curves. This delayed ionization is in concurrence with the inference from the increase in temperature evolution in Section 4.3, as it affirms the breakout of the shock from a confined dense CSM. The He II 4686 vanished in the high-cadence spectroscopic observations presented in Bostroem et al. (2023) between 7.6 d to 8.4 d indicating that the flash spectroscopy phase lasted approximately 8 days.

### v2

@@ 3.3 Imaging Polarimetry and Spectropolarimetry
We obtained spectropolarimetric observations of SN 2023ixf over five epochs on 2023 June 14, 15, 16, 18, and 19 using the CCD Imaging/Spectropolarimeter (SPOL; Schmidt et al., 1992) mounted on the 2.3 m Bok Telescope at Steward Observatory, Kitt Peak, Arizona. We used the 600 l/mm grating with a spectral coverage of 4000-7550Å and obtained a spectral resolution of 26.53 Å using a slit size of 0.4 mm (corresponding to 4.1” in the sky). In addition to this, we used a Hoya L38 blocking filter with a cut-off wavelength of 380nm. We used polarized standards (Hiltner 960; HD 155528) to calibrate the polarization angle, with the average discrepancy between the observed and expected polarization angles for multiple standards being less than 1∘ (Schmidt et al., 1992). Furthermore, the unpolarized standard (BD+28°4211) observations confirmed that SPOL exhibited low instrumental polarization, measuring at less than 0.1%. We use the same data acquisition and reduction procedure mentioned in Bilinski et al. (2018). The polarized spectra corresponding to four exposures that sample 16 waveplate positions were flat-field corrected, background subtracted, and flux calibrated using the SPOL polarization data reduction pipeline to obtain the Stokes parameters q, u, p, and polarization angle.
@@ 4 Photometric Evolution
Panchromatic light curves covering ultraviolet, optical, and near-infrared wavelengths from various facilities are shown in Figure 2, displaying the well-sampled SN evolution spanning 1.4 d – 150 d after the explosion. All the magnitudes presented in this paper were calibrated to the AB system using the transformations described in Blanton & Roweis (2007).
@@ 4.1 Rise times, decline rates and plateau-length
The light curve peaks were measured by fitting the light curves with a smooth spline, calculating the gradient () of the fit, and identifying the zero-crossing points to determine the peaks. SN 2023ixf shows a sharp rise to the peak in the ultraviolet bands - and (i.e., blue peak) with a rise time of 4.5 d. The rise time is 6 d in -band, which is faster than the prototypical rise time of 10 d in Type II SN 1999em (Leonard et al., 2002). The steeper rise to the maximum for SN 2023ixf is attributable to the interaction of the SN ejecta with CSM Förster et al. (2018). The light curves in the redder bands () seem to show a steep rise to a distinct shoulder, i.e., red shoulder post 5 d and a gradual ascent to the maximum, i.e., red peak at 16 d. The red shoulder is seen as an abrupt change of slope in the gradient of the early light curve. The presence of the red shoulder is driven by the flux excess due to the interaction of SN 2023ixf with a confined CSM since it occurs immediately after the peak in the UV wavelengths (i.e., blue peak). A similar shoulder and peak are seen in the simulated red-band light curves of SNe II with strong wind models ( 10-3 ) by Dessart et al. (2017), affirming an alternate scenario to look for signatures of early interaction in the light curves besides the rise time. The rise time for the red peak in SN 2023ixf s higher than that of SNe IIP (7.0 0.3 d) and in the ballpark of SNe IIL (13.3 0.6 d) (Gall et al., 2015).
Post-maximum, the multi-band light curves of SN 2023ixf settle onto a plateau of roughly 75 d in and other redder bands before transitioning to the radioactively-powered tail phase at 90 d. The plateau length is at the shorter end of the typical plateau length of 100 d for SNe II, hence putting SN 2023ixf amongst some of the seldom observed short plateau SNe (Hiramatsu et al., 2021b; Teja et al., 2022). We estimated the plateau decline rates of SN 2023ixf following the prescription of Anderson et al. (2014a). The -band light curve of SN 2023ixf showed an early plateau decline rate (s1) of and a late-plateau decline rate (s2) of . The late-plateau decline rate of SN 2023ixf is higher than the mean decline rate of 1.3 inferred for SNe II (Anderson et al., 2014a). The tail phase decline rate (s3) of SN 2023ixf is which is faster than the characteristic decline rate of 56Co to 56Fe (i.e., 0.98 ) indicating incomplete trapping of and -rays.
@@ 4.2 Light curve comparisons with other SNe II
The comparison of -band absolute magnitude light curve of SN 2023ixf is shown in the top panel (A) of Figure 3. Owing to the short plateau length and fast declining nature of SN 2023ixf, we derived a comparison sample consisting of normal Type IIP SNe: SN 2007od (Andrews et al., 2010), SN 2013fs (Bullivant et al., 2018), SN 2016gfy (Singh et al., 2019), 2017eaw (Szalai et al., 2019) and SN 2020tlf (Jacobson-Galán et al., 2022); short-plateau Type IIP SNe: SN 2018gj (Teja et al., 2023b), SN 2020jfo (Teja et al., 2022), and Type IIL SNe: SN 2013by (Valenti et al., 2015), SN 2014G (Terreran et al., 2016), SN 2017ahn (Tartaglia et al., 2021), SN 2018zd (Hiramatsu et al., 2021a) and SN 2020pni (Terreran et al., 2022). The peak -band luminosity of SN 2023ixf is –18.2 mag, brighter than the average peak luminosity of SNe II (i.e., –16.74 1.01 mag) inferred by Anderson et al. (2014a), which is dominated by the population of slow-declining SNe II. The peak luminosity of SN 2023ixf is similar to the SNe IIL, namely SN 2013by, SN 2014G, and SN 2017ahn, but slightly fainter than that of SN 2020pni; however, it is brighter than the majority of normal and short-plateau SNe IIP. The plateau-decline rate of SN 2023ixf is similar to that of SN 2014G ( 1.7 ) and SN 2013by ( 1.5 ). The 75 d plateau length of SN 2023ixf is similar to that of SN 2014G, SN 2013by and SN 2013fs. Overall, SN 2023ixf shows remarkable photometric resemblance in peak-luminosity, plateau decline rates, plateau length, and plateau drop to SN 2013by and SN 2014G.
@@
We further compare SN 2023ixf to SNe II with CSM interaction, followed up extensively in UV by Swift over the last two decades. The data was downloaded from the Swift Optical/Ultraviolet Supernova Archive (SOUSA777https://pbrown801.github.io/SOUSA/, Brown et al., 2014b) and the Vega-mag were transformed to AB-mag for consistency. In the exhaustive sample of SNe II observed by Swift, the and bands exhibit a rapid surge in the flux, brightening by over 3 mags in 3 days, before reaching a peak magnitude of –20 mag at 4.5 d. This prolonged brightening observed in the initial phase of the UV light curve indicates a shock breakout within a compact and dense CSM, leading to a more luminous and elongated shock breakout event (Ofek et al., 2010). Such a distinct signature in the early UV light curves has been observed in only a limited number of SNe II. The detection of a UV burst extending over 1-day in PS1-13arp by Gezari et al. (2015) was the first observation hinting at the possibility of the shock breaking out into a confined CSM. In our comparison sample, only SN 2018zd (Hiramatsu et al., 2021a), SN 2020pni (Terreran et al., 2022), and SN 2020tlf (Jacobson-Galán et al., 2022) showed a similar intensification of the early phase UV light curve.
@@
The evidence of the sharply declining UV flux was seen in the FUV spectra of SN 2023ixf obtained by Teja et al. (2023a) at 23 d, which shows a featureless spectrum owing to the line-blanketing from the iron-group elements (Bufano et al., 2009). Despite this, we still infer a decent contribution ( 10%, see Panel B in Figure 4) of UV flux to the overall bolometric light curve (i.e., a UV excess) during the mid-plateau phase ( 38 d). Our observations indicate continued interaction with CSM despite the absence of discernible signatures in the spectral sequence outlined in Section 5.2. This aligns with the inference from the theoretical models of interacting SNe II by Dessart et al. (2022) that display early interaction frequently display a surplus of UV radiation during late phases.
@@ 4.3 Bolometric Light Curve, Temperature and Radius Evolution
We computed the pseudo-bolometric light curves of SN 2023ixf using SuperBol (Nicholl, 2018). The bolometric luminosity, color temperature (), and radius () evolution of the layer of thermalization were estimated using the blackbody fits to the SED of the SN at each epoch. The missing data in certain filters over the intermediate epochs was interpolated using a low-order spline. We computed the bolometric light curve in 3 wavelength bins, i.e., UVOIR (0.16 – 2.35 m), UVO (0.16 – 0.85 m) and OIR (0.38 – 2.35 m) and is shown in Panel A of Figure 4. For the integration, the blue boundary for the UV and optical wavelengths is defined by the blue edge of the Swift band and the band, respectively. The red boundary for the optical and infrared wavelengths is defined by the red edge of the band and the band, respectively. Henceforth, we refer to the UVOIR pseudo-bolometric light curve as the bolometric light curve.
As discussed by Zimmerman et al. (2024), the color temperature and radius evolution shown in Panel C of Figure 4, shows a steep increase in the temperature from around 14,000 K to 35,000 K over a duration of 2.2 0.1 d and a relatively-flat radius evolution at (2.0 0.2) cm (13 1 AU). The extended heating at a near-flat radius, in addition to the prolonged brightening in the UV flux, as discussed in Section 4.2 are signs of further heating upon interaction with the dense CSM. From the photometric observations at 1.1 d, the bolometric luminosity, = (1.28 0.11) , and effective temperature of (14 2) rises ten folds and two folds, respectively, in a span of just 1 d. However, due to the ensued heating, the bolometric light curve of SN 2023ixf flux peaked later on 4.5 d with a luminosity of 2.5 0.3 .
The temporal evolution of UV and NIR fraction of the pseudo-bolometric luminosity is shown in Panel B of Figure 4. The contribution of the UV flux (0.16 – 0.38 m) to the bolometric flux stays at roughly about 85% even until the bolometric maximum (same as the UV peak). If we ignore the early UV data for SN 2023ixf, the OIR bolometric light curve underestimates the bolometric luminosity by an order of magnitude, emphasizing the importance of UV observations of infant CCSNe and its importance to detailed hydrodynamical modeling. During the early nebular phase, the NIR flux contributes almost 60% to the bolometric luminosity whereas the UV flux contributes roughly 5%, highlighting the importance of NIR observations in determining 56-Ni mass in SNe II.
@@
Upon recombination of all the hydrogen in the outer envelope, SNe II post the plateau phase transitions to the tail phase powered by the decay of , which thermalizes the SN ejecta through the emission of -rays and . We computed the 56Ni-mass using the relation postulated by Hamuy (2003). The mean tail luminosity of SN 2023ixf around 145 d yields a mass of . Upon comparison with SN 1987A at a similar phase which synthesized a -mass of 0.075 0.005 (Turatto et al., 1998), a -mass of is estimated for SN 2023ixf. The above techniques assumed complete trapping of -rays by the SN ejecta. However, that is not always true, especially for short-plateau SNe due to a thinner hydrogen envelope (Teja et al., 2022) and may lead to underestimation of the yield. Hence, we modeled the late-phase UVOIR bolometric light curve of SN 2023ixf spanning 100 – 150 d using the analytical formulation from Yuan et al. (2016). We derived a total -mass of 0.059 0.001 and a characteristic -ray trapping timescale of 220 3 d, indicating a short-lived trapping and a shallower envelope. It is important to note that our parameter estimates are based on data spanning up to 150 days, which may result in a slight overestimation of the 56Ni mass. Therefore, we consider this an upper limit, as there are likely additional contributions to the light curve during the tail phase from CSM interaction, as discussed in Section 4.2.
@@ 5.1 Flash Spectroscopy Phase 8 d
Earlier spectroscopic studies on SN 2023ixf (Yamanaka et al., 2023; Jacobson-Galan et al., 2023; Smith et al., 2023; Bostroem et al., 2023; Teja et al., 2023a; Hiramatsu et al., 2023; Zimmerman et al., 2024; Zhang et al., 2023) showed a prominently blue continuum coupled with highly ionized features of C IV, N IV, C III, N III and He II along with Balmer features beginning 1.1 d from the estimated date of explosion. The features suggest flash-ionization and recombination of the confined dense CSM engulfing the progenitor of SN 2023ixf. Notably, these studies observed an increase in the relative line strengths of the highly ionized species such as N IV and C IV in contrast to N III and C III, respectively, implying a delayed ionization of the CSM. The increase in line fluxes of N IV, C IV and He II noticed in the spectral sequence of SN 2023ixf spanned 4 d from the explosion (Smith et al., 2023), which roughly corresponds to the peak of UV light curves. This delayed ionization is in concurrence with the inference from the increase in temperature evolution in Section 4.3, as it affirms the breakout of the shock from a confined dense CSM. The He II 4686 vanished in the high-cadence spectroscopic observations presented in Bostroem et al. (2023) between 7.6 d to 8.4 d indicating that the flash spectroscopy phase lasted approximately 8 days.